6, Solve the inequalitThe catheti of a triangle has length 4m. If the hypotenuse of the triangle is 9m, that is the projection of this catheti on the hypotenuse.

A.     16/9

B.     4/3

C.     4y tan(x+ ∏/3) ≥ √3 


1
Expert's answer
2021-09-13T00:26:09-0400

Right triangle "ABC"


"BC=4\ m,AB=9\ m""\cos \theta=\dfrac{BC}{AB}=\dfrac{4\ m}{9\ m}=\dfrac{4}{9}"


"BD=" projection of "BC" on "AB"


"BD=BC\cos\theta=(4\ m)(\dfrac{4}{9})=\dfrac{16}{9}\ m"

The projection of this catheti on the hypotenuse is

A.     16/9